### v1

# Firstname1 Lastname1    Firstname2 Lastname2    Firstname3 Lastname3    Firstname4 Lastname4    Firstname5 Lastname5    Firstname6 Lastname6    Firstname7 Lastname7    Firstname8 Lastname8    Firstname8 Lastname8

Abstract
Transformers excel at discovering patterns in sequential data, but understanding their limits and how they learn remains a crucial topic. In this paper, we investigate the ability of Transformers to learn pseudo-random number sequences from linear congruential generators (LCGs), defined by the recurrence relation $x_{n+1}=ax_{n}+c\;\mathrm{mod}\;m$. Our analysis reveals that with sufficient architectural depth, context length, and training data variety, Transformers can predict LCG sequences with unseen moduli ($m$) and parameters ($a,c$). We demonstrate successful learning up to $m=2^{32}$, and we perform systematic analyses of the effect of dataset complexity, architecture size, and context lengths. We uncover emergent structures in the embedding layers and attention heads, which we use to reverse engineer the underlying algorithms used by the Transformer. We first consider training and testing on a single modulus, which reveals that models learn to factorize the modulus and subsequently utilize digit-wise number representations to make sequential predictions. When generalizing to unseen moduli, the models employ a greedy approach to estimate the unknown modulus from the context and then use prime factorizations to generate predictions. In the latter case, we observe a transition at a critical depth $d=3$, below which the Transformer is incapable of learning. We also find that the number of in-context sequence elements needed to reach high accuracy scales sublinearly with the period.
Machine Learning, ICML
1 Introduction
Transformer-based language models have proven to be extremely powerful sequence generative models. With copious amounts of training data and computational resources, they can identify and learn complex patterns from training corpora, resulting in numerous remarkable capabilities (attn_is_all_you_need; dosovitskiy2021an). Recent research has demonstrated that these models, when provided with sufficient context and inference compute, can acquire new patterns and capabilities without additional training through techniques such as in-context learning (radford2019language) and chain-of-thought reasoning (wei2023chain). While these models have achieved unprecedented success, understanding what underlying patterns are learned and how they learn them remains a significant challenge.
Pseudo-Random Number Generators (PRNGs) represent an interesting test case for exploring these challenges. These algorithms, which are fundamental to modern cryptography and computer science, are designed to produce outputs that pass statistical tests for randomness, but nevertheless arise from mathematical patterns that could potentially be learned by sufficiently powerful sequence models.
This intersection between Transformer models’ pattern-learning capabilities and the structured nature of PRNG outputs raises intriguing questions about both the capabilities and limitations of these models. Can Transformers learn to predict PRNG outputs given sufficient training data, model capacity and context? If so, what implications does this have for our understanding of both Transformer architectures and PRNGs? Do the Transformers learn the underlying generating algorithm or merely detect shortcuts and spurious patterns? What effect do model capacity, data variety, training methodologies and context length have on the capabilities of Transformers?
This work aims to answer these questions by focusing on learning sequences obtained from linear congruential generators (LCGs) using GPT-style autoregressive Transformers. We demonstrate how Transformers can successfully learn LCGs and training with moduli up to $m=2^{32}$. We perform interpretability analyses, uncovering emergent structures in the embedding layers, attention heads, and underlying algorithms that the Transformer uses to learn the sequences. We also perform several systematic scaling analyses to understand the effect of architecture and sequence complexity on model performance and in-context learning ability.
1.1 Related works
Our study on the learnability of PRNGs for Transformers touches on several modern and classic topics.
Interpretability and Modular Arithmetic A growing body of work uncovers the circuits, algorithms, and emergent structures learned by Transformers (sharkey2025interp; olsson2022context; Ahn2023gradient; vonoswald2023Transformers; akyurek2023what; hendel2023incontext; liu2024incontextvector). A particularly fruitful setting involves simple problems in modular arithmetic (power2022grokking; gromov2022grokking; nanda2023progress; zhong2023clock; doshi2024to; he2024learning). Our work adds to this by reverse-engineering the underlying algorithms and uncovering emergent structures in learning pseudo-random number sequences.
Cracking PRNGs: There is a classic duality between cryptography and learning theory (rivest1991cryptography), and cracking PRNGs is an important topic in cryptography. Nevertheless, deep learning-based attacks have received limited attention in the post-Transformer era. amigo2021 demonstrated that a fully-connected neural network can predict the outputs of a modified LCG with fixed (irrational) parameters $(a,c,m)=(1,\pi,1)$. In comparison, we systematically analyze the harder cases of unseen parameters using Transformers, reverse-engineer the learned algorithms, and study effects of scale and complexity.
Context-Free Grammar: LCG can also be viewed as a formal language (Type-3 regular grammar) lying within the Chomsky hierarchy (chomsky1956three). Formal languages provide an interesting setting for synthetic datasets that can be used to understand the properties of neural networks in controlled settings (deletang2023chomsky; allenzhu2024physicslanguagemodels1; cagnetta2024deep; cagnetta2024towards).
Chaotic time-series: A major application of neural networks is predicting time-series for chaotic dynamics, such as weather prediction (lam2023learning) and financial modeling. PRNGs provide an analog of such dynamics in the discrete setting.
1.2 Linear Congruential Generators
LCG is a simple PRNG that generates the next number in a sequence $\{x_{n}\}_{n=0}^{N}$ according to the linear map:
$x_{n+1}=(ax_{n}+c)\mod m,$ (1)
where $m>0$ is the modulus, $0<a<m$ is the multiplier and $0\leq c<m$ is referred to as the increment. An LCG map is uniquely defined by the choice of $m,a,c$ and the initial seed $x_{0}$. An important quantity that determines the complexity of an LCG sequence is its period: $1\leq\mathcal{T}_{m}(a,c)\leq m$. As we will show in the following sections, the period of a sequence plays a major role in the difficulty of prediction with Transformers. According to the Hull-Dobell Theorem (hull-dobell), the period $\mathcal{T}_{m}(a,c)=m$ if and only if the values of $a$ and $c$ satisfy the following criteria:
  1. I.
    $m$ and $c$ are coprime
  2. II.
    $a-1$ is divisible by all prime factors of $m$
  3. III
    $a-1$ is divisible by $4$ if $m$ is divisible by $4$.
We will evaluate all our models exclusively on sequences that obey the criteria of this theorem.
LCGs are widely utilized for their speed and simplicity, often forming the core of more complex PRNGs like PCG-64, which is used in NumPy. LCGs perform poorly at small bit sizes but improve rapidly with larger state sizes. For instance, an LCG with 88 bits of state can pass the stringent BigCrush randomness test (ONeill2014PCGA).
2 Training Setup
2.1 Dataset Generation and Evaluation
We investigate two training paradigms: Fixed Modulus (FM) and Generalization to Unseen Modulus (UM).
Fixed Modulus (FM): We first apply the Hull–Dobell Theorem to determine possible values of $(a,c)$ that maximize the sequence period for a given modulus $m$. From these possible values, we randomly sample $64$ values of each $a$ and $c$ to generate the test dataset. To generate the training dataset, we exclude these choices of $a,c$ and uniformly sample 100k LCG sequences of length $L$ (context length), with different $a,c$ and $x_{0}$ (regardless of their period).
Generalization to Unseen Modulus (UM): In this more challenging paradigm, we first select test modulus $m_{\text{test}}$ that are excluded from training. The training dataset incorporates $n_{m}$ different modulus uniformly sampled from the range $[L,m_{\text{max}}]$, where $m_{\text{max}}=\lfloor 1.2\cdot m_{\text{test}}\rfloor$. For each modulus $m$, we uniformly select $n_{a}$ multipliers ($0<a<m$) and $n_{c}$ increments ($0\leq c<m$), generating sequences of length $L$ with random initial values $x_{0}$. This results in a total of $N=n_{m}\times n_{a}\times n_{c}$ training sequences. To construct the test dataset for each $m_{\text{test}}$, we randomly sample 64 values each of $a$ and $c$ that generate maximum-period sequences.
In both paradigms, test accuracy is measured by averaging over all sequences generated from the excluded values of $a$ and $c$ and 128 random initial values $x_{0}$ for each $(a,c)$ pair.
The above settings are used in Sections 3 and 4. In order to achieve better performance, we used a larger and slightly higher-quality dataset in Section 5, which we detail later.
2.2 Tokenization, Architecture and Optimizer
We employ two different settings throughout this paper. (i) In Sections 3 and 4, we tokenize each number as a unique token, with dictionary size $m$ for the FM case and $m_{\mathrm{test}}$ for the UM case. We then use GPT-style Transformers with learnable positional embeddings and weight tying (press2017tying). (ii) In Section 5, we tokenize each number using byte-tokenization with various base choices and apply abacus positional embeddings (mcleish2024abacus)
Model architecture is characterized by number of blocks (depth), embedding dimension ($d_{\mathrm{model}}$), number of attention heads ($n_{\mathrm{heads}}$). Models are trained with AdamW optimizer.
3 Training Results
Figure 1: Training phase diagram and dynamics. (a) Test accuracy for last token of models trained with $m_{\mathrm{test}}=4096$, evaluated across various depths and $n_{\mathrm{head}}$ values, showing a minimum depth requirement of $3$ for generalization; (b) Training curves for last token of a model with depth $6$ and $n_{\mathrm{head}}=4$ ($m_{\mathrm{test}}=2048$), showing that test accuracy for last token “groks” simultaneously with performance on sequences with period longer than context length $L=256$, indicating delayed discovery of underlying rules.
In this section, we focus on the more sophisticated behavior exhibited by models trained with UM data, leaving the discussion of FM for Appendix A. We begin by investigating the minimal model that can solve the task. In Figure 1 (a), we present the test accuracy for the last token of models trained on a dataset with $m_{\mathrm{test}}=4096$, varying both model depth and number of attention heads ($n_{\mathrm{heads}}$) while keeping head dimensions constant. Our results indicate that a minimum depth of $3$ is necessary for good generalization performance, with a sharp transition upon lowering the depth. Further analysis across multiple $m_{\mathrm{test}}$ values (see Appendix B) confirms that this minimal depth requirement is universal.
We then carefully examine the training dynamics of a model with $n_{\mathrm{head}}=4$ and depth $6$ in Figure 1 (b). We categorize the training sequences into two groups: i) sequences with periods shorter than the context length, which can be solved through simple copying, and ii) sequences with periods longer than the context length, which require the model to deduce underlying rules for prediction. Our analysis reveals that the model first acquires copying ability for group i) in the early stages of training, and later “groks” the solution for group ii). Notably, the model’s ability to generalize to test $m_{\mathrm{test}}$ emerges simultaneously with this grokking phenomenon. These results demonstrate that the model develops different capabilities at distinct stages of training, with generalization ability emerging only after the model learns the underlying rules through solving the more challenging sequences. In Appendix C, we present an ablation study where models are trained exclusively on either short-period or long-period sequences. Our findings indicate that training exclusively on long-period sequences enables model generalization and eliminates the grokking phenomenon.
4 Interpreting how Transformers predict PRNGs
Interpreting the underlying algorithms implemented by the models is essential for understanding the models’ training dynamics and scaling capabilities. We uncover these algorithms for both FM and UM cases. While certain details of the algorithms differ in the two cases, they share some common properties originating from the underlying LCG structures.
We first discuss properties of LCG sequences that will be useful to interpret model behaviors.
4.1 Mixed Radix Representations
Figure 2: Bit-wise periods in an example LCG sequence generated with $m=2048,a=293,c=1033$, which follows the Hull-Dobell theorem. In binary representation $w$-th lowest bit has a period of $2^{w}$, for $w\in\{1,\dots 11\}$. Writing a new sequence by skipping every 2nd step ($r=2$) reduces the periods of all the bits by a factor of $2$, rendering the lowest bit constant. $r=2^{k}$ reduces bit-wise periods by a factor of $2^{k}$, with last $k$ digits constant.
Consider an LCG sequence with modulus $m=2048=2^{11}$. Each number in this sequence can be represented as an 11-digit binary number:
[…]
where $\{\alpha_{2^{1}},\dots,\alpha_{2^{11}}\}$ are the binary digits (bits), with $\alpha_{2^{j}}\in\{0,1\}$.
In this binary representation, each digit (bit) has its own period along the LCG sequence. As shown in Figure 2, for a sequence of period $\mathcal{T}_{m}=m$, the $w^{th}$ lowest digit has a period of $2^{w}$ along the sequence. Thus, lower (higher) digits have smaller (larger) period along LCG sequences.
While the binary representation is natural for moduli of the form $m=2^{w}$, it is possible to generalize this representation to the case of composite moduli. Consider sequences with composite modulus $m$, which has a prime factorization $m=p_{1}^{w_{1}}p_{2}^{w_{2}}\cdots p_{i}^{w_{i}}$. In this case, the numbers in the sequence can be represented uniquely in a mixed radix form:
$n=\sum_{w=1}^{w_{1}}\alpha_{p_{1}^{w}}\,p_{1}^{w}+\sum_{w=1}^{w_{2}}\alpha_{p_{2}^{w}}\,p_{2}^{w}+\cdots+\sum_{w=1}^{w_{i}}\alpha_{p_{i}^{w}}\,p_{i}^{w}\;;$ (3)
where the digits $\alpha_{p_{j}^{w}}\in\{0,1,\dots,p_{j}-1\}$ are the generalization of bits. When $\mathcal{T}_{m}=m$, the period of each digit $\alpha_{p_{j}^{w}}$ has a period of $p_{j}^{w}$. We will see later that trained Transformers have emergent structures that find such digit-wise representations and utilize them to make systematic predictions from the context. The per-digit period will play an important role in the prediction accuracy of that digit.
Consider the $r$-step iteration of Equation 1:
$x_{n+r}=a^{r}x_{n}+\sum_{i=1}^{r}a^{i-1}c\mod m\;.$ (4)
In this new sequence, the period of each digit $\alpha_{p_{j}^{w}}$ reduces from $p_{j}^{w}$ to $p_{j}^{w}/\gcd(r,p_{j}^{w})$. Consequently, some of the higher digits become relatively simpler to predict due to reduced periods; and some lower digits become trivial to predict due to being constant. We show an example of this for $m=2048$ in Figure 2 (top panel). The digit-wise periods in the new sequence with $r=2$ reduce by a factor of 2, while the last digit is simply constant. Higher values of $r=2^{k}$ will lead to even further simplifications of the sequence. We will see in the later subsections that the Transformer has emergent structure that leverages this simplification to accurately predict LCG sequences.
For general $m$, LCG sequences simplify when $r=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots$. This simplification from period-reduction is only visible when we represent the numbers in the sequence in their mixed radix representations corresponding to $m$. A different representation would only partially (or not at all) reduce the per-digit periods.
4.2 Interpretability: Fixed Modulus
Qualitative Algorithm (fixed modulus): i. Find mixed-radix representations of inputs from the learned prime factorization of $m$ ii. Look back $r=p_{j}^{k}$ steps in the context and copy the lowest $k$ digits, for different prime factors of $m$ iii. Using these $r$-step sequences, predict the higher digits of the simplified sequence
We now discuss the algorithm implemented by Transformers trained on LCG with a fixed modulus. We will show that, despite being provided integers as inputs, the model develops emergent structures that create and leverage the mixed-radix representation of the inputs. We will focus on the setup with a 1-layer, single-head Transformer trained on LCG with $m=2048$. Similar results for composite moduli (e.g. $m=7776$) and different model sizes are presented in Section D.1.
Figure 3: Test accuracy of trained model ($m=2048,l=1,n_{head}=1,d_{model}=768$), averaged over $a,c$ and initial seeds. (a) Test accuracy w.r.t. token positions. A ladder-like structure appears, with jumps occurring at $2^{k}$-th positions. (b) We represent numbers as an eleven-digit binary number ($2048=2^{11}$) and compute the per-digit test accuracy of model predictions.
We begin by analyzing the average accuracy of a trained Transformer model as a function of token position along the context, shown in Figure 3 (a). The accuracy exhibits a ladder-like structure, with each jump occurring exactly at the $2^{k}$-th token position. These successive transitions can be explained using binary representations and $r$-step recurrence (Equation 4). Specifically, once the token position exceeds $2^{k}$, the model can look back in the context and implement $r=2^{k}$-step iteration. This is utilized for (i) copying the lowest $k$ bits, since they become constant as well as (ii) simplifying the higher bits, since their periods get reduced by a factor $2^{k}$. We note that the accuracy trend remains unchanged across different choices of $a$ and $x_{0}$ (see Section D.1 Figure 18 ).
Next, we compute the per-digit accuracy by converting both model predictions and ground truth labels to their binary representations according Equation 2. In Figure 3(b), we observe that the model can predict more digits correctly with more context, with the transitions occurring at $2^{w}$-th token positions. This is a direct result of the sequences becoming increasingly simplified as the model can look farther back in the context and utilize $(r=2^{w})$-step iterations. Since this simplifications occurs in the form of digit-wise periods, Figure 3(b) serves as direct evidence that the model is internally developing and utilizing binary representations. The sequential learning of digits also explains the ladder-like structure of the accuracy in Figure 3(a). We find that the overall average accuracy (Figure 3(a)) multiplicatively depends on the per-digit accuracies (Figure 3(b)) (see Section D.1 Figure 15(b)) [^1] This holds if the per-digit accuracies are independent, which we confirm empirically.
[…]
Figure 4: Fixed modulus interpretability ($m=2048$, depth=1, $n_{heads}=1,d_{model}=768,a=1589,c=629$). (a1) 1st principal component groups the numbers mod 2. (a2) 2nd and 3rd principal components group the numbers mod 4. (b) Embedding vectors of different numbers exhibit high cosine similarity when they are spaced $2^{k}$ distance apart – with the similarity increasing with $k$. (c) Each query attends most strongly to the tokens $2^{k}$ and $2^{k-1}$ distance backward, for highest possible value of $k$, enabling copying of $k$ lower bits. (d) post-ReLU MLP activations at token position 129 with respect to different target numbers (i.e. number at the next token position). Each neuron gets activated only for specific target numbers, with a preiodic pattern.
Next, we investigate how various components of the model implement the algorithm outlined earlier this section.
Step i: We begin by analyzing how the model performs prime factorization to develop the binary representations (mixed-radix for general $m$). We conduct Principal Components Analysis (PCA) of the embedding matrix. Figure 4(a,b) shows the projections of all numbers $n\in\{0,\dots 2047\}$ along the first principle component of the embedding. We observe that the model groups the numbers into modulo $2$ clusters along the first principal component. Similarly, the 2nd and 3rd principal components group the numbers into modulo $4$ clusters. In general, we find principal directions that group the numbers into modulo $2^{w}$ clusters (see Section D.1 Figure 19). By clustering the numbers according to their remainder modulo different prime-powers, these principle directions naturally encode the digit-wise representations of the inputs. In Figure 4(c), we check the cosine similarity between the embedding vectors of different numbers. We see that the more digits two numbers share in the binary representation, the higer the cosine similarity of their embedding. This is a consequence of these numbers having similar components along principle directions corresponding to those digits.
Step ii, iii: To validate that the model looks back $r=2^{k}$ steps in the context, we examine the attention weights in Figure 4(d). In order to predict the token at position $t$, the model attends to the positions $t-{2^{k}}$. This allows the model to copy the lowest $k$ bits and simplify the prediction of higher bits. We observe that the model attends most strongly to the positions $t-{2^{k}}$ and $t-{2^{k-1}}$ for the largest possible value of $k$. This corresponds to the two brightest lines in Figure 4(d) that are ${2^{k-1}}$ distance apart. By comparing the binary representations of these two numbers, the model copies the lower $k$ bits to make predictions.
In addition to the two bright lines in the attention score, we observe multiple faint lines which are $2^{k^{\prime}}$ apart, for $k^{\prime}<k-1$. The model combines the information from all these tokens to predict the higher bits. To verify that the two bright lines at positions $t-{2^{w}}$ and $t-{2^{w-1}}$ copy lower bits, and the other faint lines are important for predicting higer bits, we perform the following ablation. For each query, we mask out all attention scores except for the top two key positions. We then measure the performance of this ablated model in Section D.1, Figure 15(d). In this case, we observe that the model retains the ability to copy the $k$ lowest bits, but loses the ability to predict higher bits.
After the attention layer gathers the information about previous tokens, MLPs implement the remaining computational processing to make predictions. In Figure 4 (d), we plot the post-ReLU MLP activations at fixed token positions w.r.t. the number appearing at that position. We find that each neuron only fires when a specific set of numbers appear at those tokens – often exhibiting periodic structure. We conclude that the first fully connected layer of the MLP determines which neurons to fire based on the context. The second fully connected layer aggregates the non-zero activations to form the final prediction.
4.3 Interpretability: Generalization to Unseen Modulus
Qualitative Algorithm (unseen modulus): i. Encode information about many possible prime factorizations ii. Use the largest number in the context to estimate the modulus iii. Implement steps ii and iii from the fixed modulus algorithm with a different realization
Unlike the FM case, the training set for UM is generated using many different moduli $m$. Since each modulus has its own mixed-radix representation, which is incompatible with the other moduli, the model needs to implement a different, more general algorithm to solve UM tasks.
We analyze how various components of a 4-layer Transformer model implement the above algorithm to solve the UM task. The model is trained on a dataset generated with $n_{m}=n_{a}=n_{c}=128$, where $m$ values are randomly chosen from the interval $[257,2457]$. To avoid leakage between train and test sets, we specifically exclude the moduli $m_{\mathrm{test}}\in\{1800=2^{3}\cdot 3^{2}\cdot 5^{2},2048=2^{11},2352=2^{4}\cdot 3\cdot 7^{2}\}$ from the training set.
((a))
Figure 5: PCA analysis of the embedding layer. The first principal component resembles a circular structure, while the second and third dimensions categorize numbers based on their modulo $2$ and modulo $3$ values, respectively.
Step i: We begin by analyzing the embedding layer. In Figure 5(a), PCA reveals that the model forms a semi-circular structure in the first two principal components. This structure resembles the circular patterns observed in modular arithmetic tasks (power2022grokking; zhong2023clock). However, since our model is trained with multiple moduli, it cannot determine the value that should be identified with 00 to form a closed circle.
Further analysis of Figure 5 (a, b) demonstrates that the $2$nd and $3$rd principal components of the embedding layer categorize numbers according to their values modulo $2$ and $3$, respectively. Given that the modulus is not fixed in this case, there is no inherent requirement for the model to encode numbers in binary representation. In fact, as we will demonstrate next, the model forms radix representations based on the choice of modulus $m_{\mathrm{test}}$ through specialized attention heads. We believe the embedding layer’s primary focus on modulo $2$ and $3$ can be attributed to these being the most common prime factors of numbers appearing in the training set.
In the first layer, attention heads further process the embedded numbers by grouping them according to their modular values with respect to various prime factors, with each head specializing in specific prime factors. This head specialization enables the model to construct radix representations requiring different prime bases.
((a))
Figure 6: PCA analysis of attention heads specializing in different prime factorizations and their effect on per-digit accuracy. (a1, b1, c1) PCA patterns of outputs extracted from corresponding heads. The first two principal components group numbers based on their values modulo $14$, modulo $4$, and modulo $3$, respectively. (a2, b2, c2) Test accuracy for individual digits when numbers are represented according to Equation 3 for given $m_{\mathrm{test}}$. (a3, b3, c3) Test accuracy for individual digits after pruning the corresponding heads shown in (a1, b1, c1). Pruning these specific heads substantially degrades performance on their corresponding digits.
The attention heads responsible for preparing different prime factorizations are shown in Figure 6. To get the (a1, a2, a3) panel, we feed the model with sequences generated using the corresponding $m_{\mathrm{test}}$ value and randomly selected legitimate values of $a$ and $c$. Then, we extract outputs from each corresponding head, at the identical token positions for all $m_{\mathrm{test}}$ possible numbers to form a $m_{\mathrm{test}}\times d_{\mathrm{model}}$ matrix. Finally, we perform PCA analysis over this matrix and project each number’s feature vector onto the top two principal components. The analysis reveals that each head naturally categorizes numbers according to their values modulo different prime factors. This head specialization enables the subsequent model layers to select which representation to use. The prominence of small prime numbers in the top principal components likely reflects their frequent occurrence in the prime factorization of composite numbers.
To further confirm the importance of these specialized heads, we compare the model’s per-digit performance before and after pruning the corresponding heads. In Figure 6 (a2, b2, c2), we measured the model’s performance before pruning, where we measured per-digit performance by representing numbers in the natural radix representation for each test modulus. The results show that the model typically struggles with calculating higher digits for a larger prime factor. This behavior aligns with what we observed in the FM case, where computing higher digits is more demanding than copying lower digits. Different from FM case, the difficulty in calculating higher digits is partly due to the model’s inability to precisely determine the modulus $m_{\mathrm{test}}$ from the sequence, which we will elaborate more later in this section.
In Figure 6 (a3, b3, c3), we measured the model’s performance after pruning the corresponding heads. By pruning, we mean replacing all components of the outputs from the corresponding heads with a single value. This value represents the mean magnitude of the outputs, computed by averaging across all dimensions: different sequences (using $10\%$ of randomly selected training sequences), token positions, and embedding dimension. Note that zeroing out any head’s contribution in the first layer would annihilate the model’s performance. In contrast, our pruning method ensures that we remove the contribution of an attention head without changing the scale of the signal drastically. Comparison with the pre-pruning results reveals that for each $m_{\mathrm{test}}$, pruning the head responsible for a specific prime factorization dramatically reduces performance on the corresponding digits while having a relatively minor impact on others. For instance, in panel (a3), the model’s ability to compute digits related to base-$7$ is completely lost, while digits related to base-$2$ and base-$3$, though impacted, remain above random chance. Similarly, panel (b3) shows complete performance degradation. In panel (c3), the model’s ability to compute digits related to $3$ is drastically decreased, while its capacity to process digits for base-$2$ and base-$5$ digits remains intact or shows marginal degradation.
To further validate the correlation between specific attention heads and digit-wise performance, we conduct pruning experiments as shown in Figures 20 and 21. Interestingly, in Figure 20 (a, b), when we remove the attention head responsible for computing modulo $3$ representations in Figure 6 (c1) and evaluate the model’s performance for $m_{\mathrm{test}}=2048$, we observe improved performance at specific token positions for certain bits, while other bits exhibit minimal performance degradation. This behavior suggests that the model experiences internal uncertainty about which radix representation to employ, likely stemming from its difficulty in accurately determining the value of $m_{\mathrm{test}}$, as we will demonstrate in the following paragraph.
((a))
Figure 7: Attention head (layer 1, head 6) for estimating $m_{\mathrm{test}}$. (a) The query consistently attends to keys corresponding to the largest values within the given sequence; (b) the head generates features with the highest cosine similarity to $m_{\mathrm{test}}$; (c) when patching the feature with sequences generated using $m_{\mathrm{patch}}=1024$ but maintaining the same $a$ and $c$, the model often makes predictions with values of $n$ smaller than $m_{\mathrm{patch}}$.
Step ii: Next, we show how the model estimates $m_{\mathrm{test}}$ in-context in Figure 7, where we find one specific attention head in the first layer that is responsible for estimating $m_{\mathrm{test}}$. In panel (a), we visualize the attention weights for an input sequence generated with $a=5$, $c=31$, and $m_{\mathrm{test}}=2048$. We find that the head focuses exclusively on the few largest numbers within the context. In panel (b), using the same sequence, we extract output features from this specific head by zeroing out the contributions from all other heads in the first layer, and compute the cosine similarity between this extracted feature and the embedding of all numbers $n$. The head consistently maintains high cosine similarity with the few largest numbers observed within the context, representing a greedy estimation of $m_{\mathrm{test}}$.
To further confirm that this head generates features corresponding to $m_{\mathrm{test}}$, we perform a patching experiment (zhang2024actpatch), as shown in panel Figure 7 (c). Here, we generate a sequence with the same $a$ and $c$ but with a different modulus $m_{\mathrm{patch}}=1024$, then use the feature from this patching sequence to replace the feature from the same head in the forward pass of the original sequence. The results show that the model frequently predicts subsequent numbers with values smaller than $m_{\mathrm{patch}}$. We also show in Figure 22 that patching any other heads might ruin the predictions but would never lead to a similar qualitative change, which further confirms that this head is the only one that helps the model estimate $m_{\mathrm{test}}$.
Step iii: After preparing the necessary features, subsequent layers in the model implement the rest of the algorithm. As in the FM case, we observe ladder structures in digit-wise accuracy for the lower digits at early token positions. This pattern, which demonstrates the copying bias, is visible in Figure 6 (b2). However, higher digits do not exhibit such ladder structures as the FM case. This discrepancy may arise because the model cannot precisely determine $m_{\mathrm{test}}$, requiring it to dynamically select the radix representation by identifying the most coherent signal at later stages. The potential competition mentioned earlier in step i further supports this argument. Nevertheless, we believe the model implements a similar algorithm for computing the higher bits in this case, which we will elaborate on in Section D.2.
5 Scaling Up the Modulus
Figure 8: Visualization of byte tokenization and abacus embedding. Abacus embedding 1 is shared by all the bytes within the integer, while Abacus embedding 2 varies within the byte but is shared by all integers.
In this section, we investigate Transformer training upon scaling up the modulus of LCG. For this purpose, we make the following modifications to our model:
Byte Representation Inspired by the emergent radix representations observed in the previous section and seeking to avoid massive dictionary sizes for large $m$, we implement a byte-level tokenization scheme. Each integer is decomposed into a sequence of bytes, beginning with the least significant byte. The vocabulary has a size of $256$ for any $m$.
Abacus Embeddings We encode positional information using a variant of the Abacus embedding (mcleish2024abacus), as a sum of two learnable vectors. One vector encodes the position of the integer within the sequence, while the other encodes the position of each byte within the integer. In Figure 8, we show a visual representation.
Fixed Modulus For FM training, we train a 2-layer model with $d_{\mathrm{model}}=1024$ and a vocabulary size of $256$. We select training sequences via the Hull-Dobell theorem, setting $n_{a}=n_{c}=1024$ (See Section F.1 for training details). For the test dataset, we choose $a$ and $c$ that differ from those in the training set. The quality of an LCG largely depends on its multiplier, traditionally evaluated via the spectral test. In Figure 9(a), we test our model on both spectrally optimal Steele multipliers (steele2021) and arbitrary multipliers for $m=2^{32}$. While achieving $100\%$ test accuracy with equal in-context sequence lengths, the model performs consistently worse on Steele-generated sequences compared to those from arbitrary multipliers.
In Figure 9(b), a log-log plot reveals that the number of in-context examples needed for $100\%$ test accuracy scales sublinearly with modulus $m$ as $m^{\alpha}$, where $\alpha\approx 1/4$.
Figure 9: (a) Average test accuracy ($m=2^{32}$) vs number of in-context examples. (b) The number of in-context examples required to achieve 100% test accuracy (minimum of 5 runs). (See Appendix E for details)
Generalization to unseen modulus For UM case, we train a 6-layer Transformer on a dataset with $n_{m}=32,768,n_{a}=128,n_{c}=1$ ($655361024<m_{\mathrm{train}}<$ 65536), with 512 numbers in each sequence. As before, we use unseen values of $m_{test}$ and $a,c$ for the test dataset – with a focus on $m_{test}$ of the form $2^{k}$ and $3^{k}$. In Figure 10, we show how the number of in-context sequence elements is required to reach $60\%$ test accuracy scales with $m$. The averaged test accuracy of each number in the sequence is shown in Section F.2. Test performance is influenced by the tokenization base, since the tokenization base highlights the periodic structure of LCGs making it more apparent and easier for the model to leverage during training and prediction. In Figure 10 we observe that $m=2^{k}$ ($m=3^{k}$) sequences offer favourable scaling when the tokenization base is $256=2^{8}$ ($243=3^{5}$).
Figure 10: In-context eample scaling for UM: The number of in-context sequence elements required to achieve 60% test accuracy shows a sublinear dependence on the modulus $m$. The scaling depends on the compatibility between $m$ and tokenization. (a) base-$2^{8}$ tokenization, (b) base-$3^{5}$ tokenization.
6 Discussion
We have investigated Transformer training on LCG sequences, focusing on fixed modulus training as well as generalization to unseen moduli. In both cases, we have uncovered the algorithm used by the model to solve these tasks, and highlighted the model components that implement the steps of the algorithm. We have found that the model finds and utilizes prime factorizations of $m$ and mixed-radix representations of numbers to simplify the sequences and make predictions. We have provided the modified training recipe for scaling up the modulus in both FM and UM settings, and shown their scaling behaviors.
Limitations and future work: While we scale the FM setting to $m=2^{32}$ the We only scale UM setting up to $m=2^{16}$ in this work. We also leave the exploration of PRNGs that are built upon LCG, such as PCG and TCG for future works.
Impact Statement
Our work advances the understanding of how neural networks learn deterministic sequences, specifically LCGs. While this capability cannot compromise mainstream cryptographic systems, which use far more sophisticated techniques, our insights may contribute to the development of more robust cryptographic algorithms and a better understanding of neural networks’ computational capabilities.
Appendix A Fixed Modules Training Results
For the FM case, we find that a single attention head in one layer is sufficient to solve the task. In Figure 11, we present the model’s training loss and performance across training steps for $m=2048$. Notably, panels (c, f) reveal a significant disparity between training and test loss, indicating a grokking transition during the training process. Similarly, we plot in Figure 12 for similar curves for $m=776$, where all curves are qualitatively the same.
Figure 11: Test accuracy and train/test loss for $m=2048=2^{11}$, depth=1. (a,b,c) $n_{heads}=1$ (d,e,f) $n_{heads}=4$.
Figure 12: Test accuracy and train/test loss for $m=7776=2^{5}\cdot 3^{5}$, depth=1. (a,b,c) $n_{heads}=1$ (d,e,f) $n_{heads}=4$.
Appendix B Critical Depth for the Unseen Modulus Task
Figure 13: Test accuracy heatmaps of
This section analyzes the depth and embedding dimension requirements for successfully training a Transformer on the unseen modulus task. We generate $N=4\times 10^{5}$ training examples covering at least $m_{\text{eval}}/4$ moduli in the training dataset.
We varied depths $\in\{2,3,4,6,8\}$ and embedding dimensions in $d_{\text{model}}\in\{512,768,1024,1280\}$, with the head dimension fixed to $d_{\text{head}}=128$. For each depth and width, we scanned learning rates $\eta\in\{3\times 10^{-5},1\times 10^{-4},3\times 10^{-4},1\times 10^{-3}\}$ and weight decay strengths $\lambda\in\{0.01,0.1,1.0,3.0\}$ to identify the optimal hyperparameters. We report that the optimal learning rate and weight decay strength heavily vary with depth and embedding dimension and training dataset. For $m_{\text{eval}}=\{1024,4096\}$, the models were trained for $T=100,000$, while for $m_{\text{eval}}=16,384$, the models required a longer training for $T=200,000$ steps. In both cases, we used a batch size of $256$ and a linear learning rate warmup over $4,096$ steps using AdamW with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.99$.
Figure 13 shows the test accuracy heatmaps with depth and embedding dimensions as the two axes. These results demonstrate that a minimum depth of $3$ is required to learn the LCG sequence prediction task, with a marginal dependence on embedding dimension. This suggests the unseen modulus task requires a minimal computational depth of three to capture the underlying structure of LCGs.
Appendix C Training Time Interpratibility
Figure 14: (left) Comparison of the training accuracy of sequences with different periods relative to the context length $512$, (center) Accuracy when the model is only trained on sequences with period $<512$, (right) Accuracy when the model is trained on sequences with period $>512$.
In this section, we examine the order in which training sequences with different periods are learned during training. For this experiment, we consider a six-layer Transformer with $4$ heads and an embedding dimension of $768$. For this experiment, we generate sequences of length $512$ using $128$ unseen moduli and $128$ values of $a$ and $c$ each. The model is trained with Adam hyperparameters: $\eta=3\times 10^{-4}$ $\beta_{1}=0.9,\beta_{2}=0.99$ and weight decay strength $\lambda=1.0$.
Figure 14(left) compares the training accuracy of sequences with different periods relative to the context length $512$. We observe that sequences with period $<512$ are memorized early in training, while the accuracy of long-period sequences coincides with the test accuracy. Next, we perform two more experiments by training the model on datasets consisting of (1) sequences with period $<256$, and (2) sequences with period $>512$. Figure 14(center, right) show the results of these experiments. The model fails to generalize when trained only on low-period sequences while training only on long-period sequences eliminates grokking.
Appendix D More Interpretability
D.1 Fixed Modulus
In this subsection, we show extra results on the model’s behavior…
Figure 15: Test accuracy for $m=2048=2^{11}$, depth=1, $n_{heads}=1,d_{model}=768$. (a) Test accuracy averaged over $a,c$ and initial seeds. (b) Multiplication of per-digit test accuracies – matches exactly with the average test accuracy. (c) Per-digit accuracy in binary representation. (d) Per-digit accuracy of an ablated model: For each query, mask out all attention weights except for the top 2, which are $2^{k}$ and $2^{k-1}$ distance backwards in the context
Figure 16: Test accuracy for $m=2048=2^{11}$, depth=1, $n_{heads}=4,d_{model}=768$. (a) Test accuracy averaged over $a,c$ and initial seeds. (b) Multiplication of per-digit test accuracies – matches exactly with the average test accuracy. (c) Per-digit accuracy in mixed radix representation.
[…]
Figure 18: Test accuracy vs token positions for $m=2048=2^{11}$, depth=1, various values of $a,c$. (a,b) $n_{heads}=1$ (c,d) $n_{heads}=4$. The accuracies for different $a$ and $c$ are exactly on top of each other.
Figure 19: Projections along Top 6 principle components of the embedding matrix, for $m=512$, depth=1, $n_{heads}=1,d_{model}=768$.
D.2 Unseen Modulus
Pruning heads corresponding to irrelevant prime factors
To further validate our findings from Figure 6 regarding the correlation between attention heads and digit-wise accuracy, we conduct additional pruning experiments. The results of these experiments are presented in Figures 20 and 21.
((a))
((b))
((c))
Figure 20: Per-digit accuracy for $m_{\mathrm{test}}=2048$ after pruning specific attention heads. (a) Results from the same model used in Section 4.3, identical to Figure 6 (b2); (b) Performance after pruning the attention head responsible for grouping numbers by their values modulo $3$, which is irrelevant for solving sequences with $m_{\mathrm{test}}$ (containing only the prime factor $2$). After pruning, the model’s performance shows marginal improvement for specific early bits at lower token positions; (c) Performance after pruning the attention head responsible for grouping numbers by their values modulo $14$. The model’s performance on $m_{\mathrm{test}}$ decreases significantly, as this head partially contributes to processing relevant prime factors. However, since the primary head responsible for binary representation remains intact, the model maintains partial functionality.
((a))
((b))
((c))
Figure 21: Per-digit accuracy for $m_{\mathrm{test}}=2352$ after pruning specific attention heads. (a) Results from the same model used in Section 4.3, identical to Figure 6 (b2); (b) Performance after pruning the attention head responsible for grouping numbers by their values modulo $3$, which is relavent for one specific digit in this case. After pruning, the model’s performance shows a clear degradation, with the strongest one happening exactly at the digit corresponding to modulo $3$ (c) Performance after pruning the attention head responsible for grouping numbers by their values modulo $2$. The model’s performance on $m_{\mathrm{test}}$ got obliterated, as there are many base-$2$ digits in the radix representation of $m_{\mathrm{test}}$ in this case.
Patching other heads
In Figure 22, we present patching experiments for heads not previously shown in Figure 7. We focus only on selected heads where patching significantly affects predictions. Notably, none of these cases exhibit qualitative changes stemming from modulus alterations, further supporting our assertion that the head shown in Figure 7 is specifically responsible for estimating $m_{\mathrm{test}}$.
((a))
((b))
((c))
((d))
Figure 22: Patching experiments following the setting of Figure 7 in the main text. None of these heads, after patching, make the model believe that the modulus is close to $m_{\mathrm{patch}}$.
Evidence for Step iii
In Figure 23, we present attention patterns and token distance statistics for a selected attention head in later layers, with the model the same as the one used in Section 4.3. The token distances are measured by the spacing between keys with top-$4$ attention weights for a given query. Our analysis reveals that the model develops a head capable of dynamically adjusting their lookback distance for computations based on $m_{\mathrm{test}}$, which we interpret as evidence for step iii of the algorithm proposed in Section 4.3.
Figure 23: Attention patterns and token distance statistics for layer 2, head 3 of the model analyzed in Section 4.3. (a1, b1) Results for a sequence with $m_{\mathrm{test}}=2048$, $a=5$, and $c=31$. The statistics reveal that the model consistently looks back at distances that are multiples of $4$, which divides $m_{\mathrm{test}}$ and enables the correct copy behavior. (a2, b2) Results for a sequence with $m_{\mathrm{test}}=2352$, $a=85$, and $c=5$. In contrast to panels (a1, b1), the same attention head now consistently looks back at distances that are multiples of $14$, which allows the model to copy lower digits from the context. This adaptive behavior demonstrates that the model has acquired the ability to dynamically adjust its lookback distance by $r$ iterations to copy the lower digits and leave the higher digits to the later layers for computation.
Appendix E Training Variance
The model may converge to different solutions depending on the random seed used for initialization and the random seed for dataset generation. LABEL:{fig:med_fixp_appendix} shows the median value across five runs, with the shaded region representing the range between the minimum and maximum values. For larger moduli, not all models successfully find a solution that achieves $100\%$ test accuracy.
Figure 24: The median number of in-context sequence elements required for $100\%$ test accuracy across five runs. The min-max range is shown as a shaded region.
Appendix F Scaling Up the Modulus
F.1 Fixed Modulus
For each modulus $m=2^{k}$, where $k\in[16,32]$, we train a 2-layer GPT model with an embedding dimension of 1024 and a vocabulary size 256. The train set consists of $n_{a}=1024$ multipliers and $n_{c}=1024$ increments, selected via the Hull-Dobell theorem. One LCG sequence of length 512 is included in the train set for each (a, c) pair, resulting in a total training set size of $n_{a}\times n_{c}=1,048,576$. For each modulus, the model is trained for 200,000 steps with a batch size of 512. The context length is 511×511$\times$511 × the digit length of $p$ in the byte representation. For example, when $m=2^{28}$, the digit length in byte representation is 4, while for $m=2^{23}$, the digit length is 3.
F.2 Unseen Modulus
We train a 6-layer GPT model on a dataset that comprises $n_{m}=32,768$ moduli, with $n_{a}=128$ training multipliers and $n_{c}=1$ training increment per modulus. This results in a total of $n_{m}\times n_{a}\times n_{c}=4,194,304$ sequences of length 512 in the training set. The moduli are sampled from $m_{train}\in\{x\in\mathbb{Z}\mid 1024\leq x<65536\}$ in Fig. 25a, from $m_{train}\in\{x\in\mathbb{Z}\mid 1024\leq x<59049\}$. Multipliers are selected based on the Hull-Dobell theorem when sufficient qualifying $a$ values are available; otherwise, random $a$ values are used to ensure $128$ multipliers for each modulus. The models are trained on 4 million sequences for 2 million steps with a batch size of $128$. Because LCGs are typically defined for moduli that are powers of prime numbers, the model is tested on moduli that are powers of the primes 2, 3, 5, and 7. The test set consists of 512 unseen $a$ values and 64 unseen $c$ values for each test modulus.
Test performance is influenced by the tokenization base, exhibiting a bias toward moduli that share the same base as the tokenization method. For instance, in Fig. 25 (a) when using a byte-level representation, the model achieves better performance on moduli $p=2^{k}$ compared to $p=3^{k}$, $5^{k}$, or $7^{k}$. As contrasted with Fig. 25 (b) where the tokenization base is $243=3^{5}$, the model performs better on moduli $p=3^{k}$. This behavior is likely due to the property of LCGs, where for moduli that are powers of a prime $b$, the lowest $k$-th digit exhibits a period of $b^{k}$. Tokenization in such a base highlights this periodic structure, making it more apparent and easier for the model to leverage during training and prediction.
((a))
((b))
Figure 25: Test accuracy vs Number index. In (a), the moduli 2048 and 16384 (blue curves) have the same root 2 as the tokenization base 256. The model performs better on these two moduli. In (b), the moduli 2178 and 19683 (orange curves) have the same root 3 as the tokenization base 243. The model performs better on these two moduli.